OpenStudy (anonymous):

Solving indefinite integrals with substitution: What do I do when there is a term outside of my u group that is a higher power than my du. Can I square or cube my du? And if so what happens after that?

OpenStudy (anonymous):

Question is \[\int\limits 18x^3(3x^2+1)^{1/2} dx\]

OpenStudy (kainui):

Yes, du doesn't really work like that, you have some misconception that I'll try to help you figure out.

OpenStudy (kainui):

Ahhh I see what's going on. You can't do this substitution, you need to find another way.

OpenStudy (anonymous):

Oh really? Greeeeaat. But thank you for the help!

OpenStudy (kainui):

You should try using 3x^2+1=u. I think that'll work.

OpenStudy (kainui):

It does. =D

OpenStudy (kainui):

You have to be clever though, so just think about it for a second and hopefully you notice something to make this substitution work for you. ;D

OpenStudy (anonymous):

What do I need to do to eliminate the 18x^3?

OpenStudy (kainui):

Alright so u=3x^2-1 right? What is dx? I'll help walk you through it.

OpenStudy (australopithecus):

u = 3x^2 + 1 du/dx u = 3x^2 + 1 du/dx = 6x du = 6xdx du/6x = dx do the substitution, and you will see

OpenStudy (australopithecus):

but kainul can explain it to you I guess

OpenStudy (kainui):

Yeah, it's whatever lol, it's a pretty fun integral honestly.

OpenStudy (anonymous):

We have differing ideas of fun. My prof just taught us substitution in 20 mins today and our final is next week heavily based on it.

OpenStudy (kainui):

Well if you take cal 2, u sub will be one of your best friends/most used tool in math toolkit next to trig sub and integration by parts. =D

OpenStudy (australopithecus):

I could totally just see myself solving this with a substitution then just doing by parts twice

OpenStudy (australopithecus):

but I'm rusty

OpenStudy (kainui):

By parts? I don't think you need more than u-substitution here.

OpenStudy (australopithecus):

yeah probably not

OpenStudy (kainui):

Ifstretc have you figured it out yet or still working on it and need help?

OpenStudy (anonymous):

So I may have done it wrong, but putting dx=du/6x into the equation (as well as the u, thats easy) i end up with 3x^2(u)^(1/2)du

OpenStudy (kainui):

Exactly! Now look back at what you chose for your u. Notice how it has x^2 in there?

OpenStudy (anonymous):

oooooooh

OpenStudy (kainui):

Yeah, see what I mean by fun?

OpenStudy (anonymous):

But +1?

OpenStudy (kainui):

Subtract 1 from each side.

OpenStudy (kainui):

Don't forget you still have the power of algebra by your side in the war against calculus.

OpenStudy (australopithecus):

Wow I need sleep I cant believe I didnt see that

OpenStudy (kainui):

\[\int\limits_{}^{}(u-1)u^{1/2}du\]

OpenStudy (australopithecus):

(u - 1) = 3x^2

OpenStudy (anonymous):

Okay so now i have that. so take the integral of that...

OpenStudy (australopithecus):

now to solve just multiply u^(1/2) into (u-1)

OpenStudy (kainui):

Distribute out the u^(1/2) and you basically have a regular old polynomial to integrate. No problems?

OpenStudy (australopithecus):

and split the integral

OpenStudy (anonymous):

so\[u ^{3/2}-u ^{1/2}du\]

OpenStudy (australopithecus):

yeah looks good

OpenStudy (australopithecus):

now split it and take the antiderivative of each

OpenStudy (australopithecus):

and dont forget + c

OpenStudy (kainui):

And plug all your stuff back in terms of x, that's also important =P

OpenStudy (anonymous):

Okay perfect. You two both deserve a huge thank you for walking my calc challenged self through that!

OpenStudy (kainui):

Haha no problem, I think it was fun for both of us. Calculus gets trickier and more clever the further on you go. I get excited sometimes when I see hard problems now because they're so satisfying to figure out.

OpenStudy (anonymous):

Fortunately for myself this will be my only calculus class in uni. But it has been challenging and at times rewarding :)

OpenStudy (australopithecus):

this problem wasn't that hard kainui, it can be solved without really thinking using a method I stated above but it would have been long and tedious

OpenStudy (kainui):

This one is so good at misleading you straight to the 3x^2 from x^3, they're kind of jerks for that lol.

OpenStudy (australopithecus):

but yeah this was a tricky problem :)

OpenStudy (australopithecus):

wait no I'm wrong lol

OpenStudy (kainui):

I suppose, but I don't think integration by parts was taught to me when I was in cal 1, sounded like he just learned u-substitution. I can imagine the u's and dv's and stuff

OpenStudy (australopithecus):

I was taught all major integration techniques in calc 1 :\ like 50% of my exam was integration

OpenStudy (anonymous):

We've only just skimmed it in our last week of calc. There was a lot of focus on curve sketching in my course

OpenStudy (australopithecus):

Yeah I dont think my method would work actually so disregard that

OpenStudy (kainui):

Damn, that sucks. I think the point where I decided I liked calculus a lot was when we started revolving functions around axes and finding the volume. It's a lot of dividing by 0 and multiplying by infinitely small things and adding infinite numbers together that gets me into it honestly lol.

OpenStudy (anonymous):

Sounds interesting, but not for me. :)

OpenStudy (kainui):

I really have to wonder about how crazily smart the people who invented all this stuff were. It's kind of like getting a glimpse of what it's like to be a genius because you get to use their mental problem solving models. It's kind of weird and fascinating. Anyways I can talk about math all night long... So good luck, I'm going to bed.

OpenStudy (anonymous):

Thank you again!

OpenStudy (sirm3d):

you can write \[\large \int_{}{}18x^3(3x^2+1)^{1/2}dx=\int_{}{}18x^2(3x^2+1)^{1/2}(xdx)\]